How much would he charge?
As his consulting fee Mr. Grave charges $90 for an 8 hour day. How much would he charge for 3 hours?

Answer by fcabanski(1391)

Use a ratio: dollars/hours.


90/8 = x/3


Cross multiply.


8x=270


Divide by 8.


x = $33.75


Alternatively find his hourly rate.


$90/8 = $11.25 per hour.


Multiply that by the number of hours.


11.25 * 3 = $33.75.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com